A protein, that is thought to be Dystrphin (0.00100 g) is dissolved in 1.00 mL of solution. The measured osmotic pressure is 1.12 torr at 25.0°C. (a) Calculate the molar mass of the Dystrphin. (b) Do you think this protein is actually Dystrphin? Explain.
Answer #1

Use relationship;

Osmotic pressure = i × M × R ×T

Where, i = Vant's Hoff factor of solute , i=1 for organic solute .

M = molarity = moles/volume of solution in litres.

R = gas constant = 0.0821 L atm /mol K

T = temperature in Kelvin = 273+25) = 298K

Osmotic pressure = 1.12 torr = 0.001474atm ( 1torr= 0.001316atm).

Putting values for M ;

M = osmotic pressure/iRT = [0.001474atm/(1×0.0821Latm/mol K × 298K)] = 0.00006025M

a) . Moles of Dystrophin = molarity × volume in litres of solution = 0.00006025M × 0.001L = 0.0000000603 moles .

Molar mass = mass/moles = 0.001g/0.0000000603 moles = 16583.75 g/mol

Thus experimental molar mass of Dystrophin = 16583.75g/mol

b) this large molar mass indicates that this is the dystopian protein.